# Beyond Accuracy-Fairness: Stop evaluating bias mitigation methods solely on between-group metrics

### v1

@@ 3.1 Materials
We will use several real world datasets that are common in the domain of fair machine learning (Le Quy et al., 2022). The Adult Income dataset contains information extracted from the 1994 census data with as target variable whether the income of a person exceeds \$50,000 a year or not. The Compas dataset includes demographic information and criminal histories of defendants from Broward County, and is used to predict whether a defendant will re-offend within two years. The Dutch Census dataset represents aggregated groups of people in the Netherlands for the year 2001, and can be used to predict whether a person’s occupation can be categorized as a high-level (prestigious) or a low-level profession (Van der Laan, 2000). The Law admission dataset contains a Law School Admission Council (LSAC) survey conducted across 163 law schools in the United States in 1991 (Wightman, 1998) and predicts whether the student will pass the bar exam or not. The Student Performance dataset describes the achievements of students in two Portuguese schools (Cortez and Silva, 2008). The classification task is to predict whether they score above average in mathematics.
@@
Name # instances # attributes Protected attribute Protected group Target attribute Positive rate Adult 48,842 10 Gender Female High income 23.93% Compas 5,278 7 Race African-American Low risk 52.16% Dutch Census 60,420 11 Gender Female High occupation 52.39% Law admission 20,798 11 Race Non-White Pass the bar 88.97% Student Performance 649 29 Gender Male High score in mathematics 53.62%
@@ 3.2.1 Machine learning models
We train a fully connected neural network on each dataset, utilizing binary cross-entropy as the loss metric. [^2] We use a state-of-the-art machine learning model, with one hidden layer with 200 nodes. We use 50 epochs, a batch size of 128 and the Adam optimizer. We do not perform hyperparameter tuning. The choice of neural networks is driven by the necessity to accommodate one of our bias mitigation methods, Adversarial Debiasing, which exclusively operates with this type of model. Although the remaining methods are model-agnostic, we opt for consistency in our approach, employing a neural network across all bias mitigation methods to ensure comparability in our final results.
@@ 3.2.2 Bias mitigation methods
As preprocessing methods, we use Learning Fair Representations (LFR) and Disparate Impact Remover (DIR). The idea behind Learning Fair Representations (Zemel et al., 2013) is that a new representation Z is learned that removes the information correlation with the sensitive attribute, but preserves the other information about X as much as possible. Disparate Impact Remover (Feldman et al., 2015) modifies the training data to reduce the influence of sensitive attributes, but preserves rank-ordering within groups.
@@
As postprocessing methods, we use Equalized Odds Postprocessing (EQP), Reject Option Classification (ROC) and Threshold Optimation (TO). Equalized Odds Postprocessing (Hardt et al., 2016) will solve a linear program to find the probabilities with which to change output labels in order to optimize equalized odds, while Reject Option Classification (Kamiran et al., 2012) will flip the predictions the model is not confident of. [^3] Reject option classification identifies instances where the model is uncertain about its prediction and essentially ‘rejects’ making a definite decision. In this implementation, aiming to enhance fairness, the labels of these instances are flipped to satisfy a fairness criteria. Threshold Optimization (Kamiran et al., 2012) is maybe the most straightforward method of mitigating bias as it will optimizes the thresholds for both groups in isolation. [^4] For this bias mitigation method, we use the implementation available through the Fairlearn toolbox (Bird et al., 2020) as this method is not available in the AIF360 toolbox. ROC and TO are implemented to enforce Demographic Parity, while EOP enforces Equalized Odds by default.
@@ 3.3.1 Performance metrics
Most benchmark studies compare the different mitigation methods on accuracy, which measures how often the prediction label assigned by the machine learning model coincides with the true label (Chen et al., 2023; Hort et al., 2021; Krco et al., 2023).
It is commonly acknowledged that accuracy is not always the best metric to measure the performance of a machine learning model, as it is for example not suitable to deal with imbalanced class distributions (as in this case a model can obtain a high accuracy by just predicting all samples as the majority class) (Chen et al., 2023; Mittelstadt et al., 2023). This has led some studies to include other metrics such as the F1-score, Precision, or Recall (Chen et al., 2023). However, another notable drawback of these performance metrics is that they measure the performance at a specific classification threshold, as they use the prediction labels of the machine learning model and not the prediction scores. We can evaluate the performance of the prediction scores by using the Area Under the ROC Curve (AUC).
AUC allows for an objective comparison across classifiers, as it is unaffected by the choice of threshold or the frequency of classes (Hand, 2009). It measures how well the prediction scores (and thus the ranking) of a machine learning model distinguishes between positive and negative cases. The formula for the AUC score:
@@
This formula means that the AUC score is equivalent to the probability that a classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance. Provost et al. (1998) previously advocated for the adoption of AUC as a standard for comparing classifiers in the broader field of machine learning. Despite this, its integration into Fair Machine Learning has been limited. Furthermore, the specific context of Fair Machine Learning presents additional justifications for focusing on the prediction scores generated by machine learning models, rather than solely on their prediction labels, due to unrealistic positive rates and the unsatisfaction of required fairness metrics.
In this study, we operate under the assumption that label bias is absent, meaning that the actual labels accurately represent the intended prediction target (Wick et al., 2019; Favier et al., 2023; Lenders and Calders, 2023). Note that if label bias was present, it would compromise the validity of both the AUC and accuracy metrics, as these measures rely on these labels for their calculations.
Lastly, we will measure the positive classification rate of the whole population. As mentioned earlier, this is important to be able to compare the different bias mitigation methods. The formula for the Positive Rate (PR):
@@ 3.3.2 Fairness metrics
Many possible fairness metrics exist, but we will report two metrics that are commonly used in the fairness domain to measure disparities between groups: Demographic parity (or statistical parity) states that the positive classification rate must be approximately the same in the protected group as in the privileged group. Statistical Parity Difference measures the difference in this positive rate between both groups. The formula for the Statistical Parity Difference (SPD):
$P(\hat{Y}=1|S=s)-P(\hat{Y}=1|S=ns)$
@@ 4.1 Whom do the bias mitigation strategies affect?
The lack of transparency in the field of fair machine learning has been acknowledged in literature (Rudin et al., 2020; Wachter et al., 2021; Goethals et al., 2023), while some have already specifically criticized the opacity in the exact effects of the different bias mitigation methods (Krco et al., 2023; Favier et al., 2023; Holstein et al., 2019). We aim to shed light on the operational dynamics of the different bias mitigation methods, by comparing the score distributions after deploying each bias mitigation method with the score distributions of the initial machine learning model. We illustrate the prediction scores from the initial ML model on the x-axis and the scores post-application of various bias mitigation methods on the y-axis. Additionally, we categorize the instances based on their affiliation with either the protected group (represented in dark blue) or the privileged group (represented in light blue). The chart is divided into four quadrants, each depicting the classification of instances as either positive or negative by the initial ML model and by each bias mitigation method. The presented results in Figure 2 are calculated using the Compas dataset. The figures for the other datasets can be found in the Appendix (Figures 5, 6, 7 and 8), with the results being consistent to the results of the Compas dataset.
@@
Figure 2: Score distributions for the Compas dataset. The x-axis represents the prediction scores of the initial ML model, while the y-axis represents the prediction score after applying each bias mitigation method.
In Figure 1(a), we observe the prediction scores of the initial ML model. By default, machine learning models use a classification threshold of $0.5$, categorizing instances above this threshold as positive and those below it as negative. This behavior remains consistent for both the protected and privileged groups.
@@
In Figures 1(b)-1(e), we can assess the relation between the prediction scores of the initial ML model and the prediction scores after applying preprocessing (LFR and DIR) or inprocessing (ADV and MFC) bias mitigation methods. Figure 1(b) reveals that LFR significantly alters the prediction scores, in a seemingly random way. Instances that receive a positive label post-LFR are notably different from those in the initial score ranking. This suggests a substantial transformation in the feature space due to the preprocessing method. We will assess how this affects the performance and fairness metrics in Section 4.2. As shown in Figures 1(c), 1(d), and 1(e), the methods DIR, ADV, and MFC also introduce significant alterations to the prediction scores, and thus the inherent ranking of instances. However, the altered prediction scores align more closely with the initial prediction scores compared to those obtained using LFR.
@@
Figure 2 also provides us with insights into the fundamental distinction among various bias mitigation methods—specifically, whether they impact the prediction scores or only the prediction labels. Unlike preprocessing and inprocessing methods, postprocessing methods do not alter the inherent ranking; instead, they adjust labels based on the prediction scores of the initial machine learning model. Each method implements a unique strategy for label flipping, all with the objective of meeting specified fairness metrics. For instance, ROC (Figure 1(f)) targets the most uncertain instances for label flipping, TO (Figure 1(h)) establishes distinct thresholds for each group, while EOP (Figure 1(g)) determines the ideal quantity of labels to flip within each group and executes these flips randomly. In the scenarios of ROC and TO, group-specific new thresholds are determined (as illustrated in Figures  1(f) and 1(h)). Regarding EOP, label flipping occurs randomly among negatively classified individuals in the protected group and positively classified individuals in the privileged group (as shown in green and red in Figure 1(g)). In this case, the initial ranking impacts the initial labels, but will have no influence on the decision regarding which labels are flipped.
@@
Our examination provides insights into the specific instances affected by each bias mitigation method, in response to the work of Krco et al. (2023). Take, for instance, the case of ROC: the instances that are flipped from a negative classification to a positive classification, are those instances that were initially classified as negative but that posses the highest prediction score according to the initial machine learning model.
@@
Table 2: Table with the similarity between the initial ranking produced by the machine learning model and the ranking produced by the model after using a bias mitigation method, measured by the Kendall-Tau statistic. The values between brackets present the similarity in ranking for the protected group and privileged group respectively. All values are between zero and one, and higher values point to more similarity in ranking.
@@
Instead of visualizing how the prediction scores of the initial ML relate to the prediction scores after each bias mitigation method, we can also calculate the overlap between the two rankings by utilizing the Kendall-Tau statistic (Kendall, 1948). The outcomes are presented in Table 2. Table 2 illustrates that LFR consistently presents the least similarity among rankings. In contrast, the use of DIR leads to a significantly higher degree of overlap in rankings across all datasets. The similarity in rankings when applying ADV and MFC falls into a more moderate category. These observations highlight the varying degrees of impact each method exerts on the prediction scores and thus the ranking of individuals. With postprocessing methods, the ranking remains identical to the ranking of the initial model, resulting in a ranking similarity of 1 (refer to Table 2), and the AUC remains consistent with the initial model (see Table 3). Our analysis reveals that across all datasets, the prediction scores and consequently the intrinsic rankings are substantially modified by every preprocessing and inprocessing bias mitigation method we evaluated. While such modifications are not necessarily negative, it is crucial to assess whether they enhance or worsen the ranking, and to what extent. The specifics of these findings will be presented in Section 4.2. We propose that for the dual objectives of fairness and accuracy optimization, preprocessing and inprocessing mitigation methods should be adopted only if they improve subgroup rankings. Should they fail to do so, utilizing the ranking produced by the original machine learning model and then applying fairness interventions post-hoc may be a more effective strategy.

### v3

@@ 4.1 Materials
We will use several real world datasets that are common in the domain of fair machine learning (Le Quy et al., 2022). The Adult Income dataset contains information extracted from the 1994 census data with as target variable whether the income of a person exceeds \$50,000 a year or not. The Compas dataset includes demographic information and criminal histories of defendants from Broward County, and is used to predict whether a defendant will re-offend within two years. The Dutch Census dataset represents aggregated groups of people in the Netherlands for the year 2001, and can be used to predict whether a person’s occupation can be categorized as a high-level (prestigious) or a low-level profession (Van der Laan, 2000). The Law admission dataset contains a Law School Admission Council (LSAC) survey conducted across 163 law schools in the United States in 1991 (Wightman, 1998) and can be used to predict whether the student will pass the bar exam or not. The Student Performance dataset describes the achievements of students in two Portuguese schools (Cortez and Silva, 2008). The classification task is to predict whether they score above average in mathematics.
@@
Name # instances # attributes Protected attribute Protected group Target attribute Base rate Adult 48,842 10 Gender Female High income 23.93% Compas 5,278 7 Race African-American Low risk 52.16% Dutch Census 60,420 11 Gender Female High occupation 52.39% Law admission 20,798 11 Race Non-White Pass the bar 88.97% Student Performance 649 29 Gender Male High score in mathematics 53.62%
@@ 4.2.1 Machine learning models
We train fully connected neural networks on each dataset, utilizing binary cross-entropy as the loss metric. [^3] We use a neural network, with one hidden layer with 200 nodes. We use 50 epochs, a batch size of 128 and the Adam optimizer. We do not perform hyperparameter tuning. Neural networks are chosen because the implementation of one the bias mitigation methods (Adversarial Debiasing) requires this. Although the remaining methods are model-agnostic, we opt for consistency in our approach, employing a neural network across all bias mitigation methods to ensure comparability in our final results.
@@ 4.2.2 Bias mitigation methods
As preprocessing methods, we use Learning Fair Representations (LFR) and Disparate Impact Remover (DIR). The idea behind Learning Fair Representations (Zemel et al., 2013) is that a new representation Z is learned that removes the information correlated with the sensitive attribute, but preserves the other information about X as much as possible. Disparate Impact Remover (Feldman et al., 2015) modifies the training data to reduce the influence of sensitive attributes, but preserves rank-ordering within groups.
@@
As postprocessing methods, we use Equalized Odds Postprocessing (EOP), Reject Option Classification (ROC) and Threshold Optimation (TO). Equalized Odds Postprocessing (Hardt et al., 2016) will solve a linear program to find the probabilities with which to change output labels in order to optimize equalized odds, while Reject Option Classification (Kamiran et al., 2012) will flip the predictions the model is not confident of. [^4] Reject option classification identifies instances where the model is uncertain about its prediction and essentially ‘rejects’ making a definite decision. In this implementation, aiming to enhance fairness, the labels of these instances are flipped to satisfy a fairness criteria. Threshold Optimization (Kamiran et al., 2012) is maybe the most straightforward method of mitigating bias as it will optimizes the thresholds for both groups in isolation. [^5] For this bias mitigation method, we use the implementation available through the Fairlearn toolbox Bird et al. (2020) as this method is not available in the AIF360 toolbox. ROC and TO are implemented to enforce Demographic Parity, while EOP enforces Equalized Odds by default.
@@ 4.3.1 Performance metrics
Most benchmark studies compare the different mitigation methods on accuracy, which measures how often the prediction label assigned by the machine learning model coincides with the true label (Chen et al., 2023; Hort et al., 2021; Krco et al., 2023). It is commonly acknowledged that accuracy is not always the best metric to measure the performance of a machine learning model, as it is for example not suitable to deal with imbalanced class distributions (as in this case a model can obtain a high accuracy by just predicting all samples as the majority class) (Chen et al., 2023; Mittelstadt et al., 2023). This has led some studies to include other metrics such as the F1-score, Precision, or Recall (Chen et al., 2023). However, another notable drawback of these performance metrics is that they measure the performance at a specific classification threshold, as they use the prediction labels of the machine learning model and not the prediction scores. We can evaluate the performance of the prediction scores by using the Area Under the ROC Curve (AUC).  [^6] Note that this ROC, which stands for Receiver Operating Characteristic is different from the Reject Option Classification, that we also shorten as ROC.
AUC allows for an objective comparison across classifiers, as it is unaffected by the choice of threshold or the frequency of classes (Hand, 2009). It measures how well the prediction scores (and thus the ranking) of a machine learning model distinguishes between positive and negative cases. The formula for the AUC score, where $S(x_{i},a_{i})$ notes the prediction score of instance $i$ where $x_{i}$ are the attributes of the individual and $a_{i}$ is their sensitive attribute:
@@
This formula means that the AUC score is equivalent to the probability that a classifier will rank a randomly chosen positive instance higher than a randomly chosen negative instance. Provost et al. (1998) previously advocated for the adoption of AUC as a standard for comparing classifiers in the broader field of machine learning. Despite this, its integration into Fair Machine Learning has been limited. [^7] One example is where Fong et al. (2021) investigate how acquiring additional features can improve the AUC of the disadvantaged group. Furthermore, the specific context of Fair Machine Learning presents additional justifications for focusing on the prediction scores generated by machine learning models, rather than solely on their prediction labels, due to unrealistic positive rates and the unsatisfaction of required fairness metrics. In this study, we operate under the assumption that label bias is absent, meaning that the actual labels accurately represent the intended prediction target (Wick et al., 2019; Favier et al., 2023; Lenders and Calders, 2023). Note that if label bias was present, it would compromise the validity of both the AUC and accuracy metrics, as these measures rely on these labels for their calculations.
Lastly, we will measure the positive decision rate (or positive classification rate) on the whole population. As mentioned earlier, this is important to be able to compare the different bias mitigation methods. We use $Y$ to denote the actual target label, and $\hat{Y}$ to denote the predicted label by the machine learning model. $A$ represents the sensitive attribute, where $a$ represents the protected group and $\neg a$ the privileged group. The formula for the Positive Decision Rate (PDR):
@@ 4.3.2 Fairness metrics
Many possible fairness metrics exist, but we will report two metrics that are commonly used in the fairness domain to measure disparities between groups: Demographic parity (or statistical parity) states that the positive decision rate must be approximately the same in the protected group as in the privileged group. Statistical Parity Difference measures the difference in this positive decision rate between both groups. The formula for the Statistical Parity Difference (SPD):
$P(\hat{Y}=1|A=a)-P(\hat{Y}=1|A=\neg a)$
@@
4.3.3 Rank correlation
We can measure the correlation between two ranked lists by using the Kendall Tau metric, denoted as $\tau$. This metric measures the similarity between the orderings of two lists by quantifying the number of pairwise disagreements between them.
The formula for the Kendall Tau coefficient ($\tau$) is defined as:
number of concordant pairsnumber of discordant pairsnumber of $\frac{\text{number of concordant pairs}-\text{number of discordant pairs}}{\text{number of pairs}}$
A concordant pair refers to a pair of observations where the order of the ranks is the same between both lists, while a discordant pair refers to a pair of observations where the order of the ranks is reversed between the two lists. A $\tau$ value of 1 indicates perfect agreement between two rankings, whereas a value of -1 indicates perfect disagreement. Values closer to zero suggest little to no correlation, implying a lack of consistency in the ranking order between the two groups being compared.
@@ 5.1 Whom do the bias mitigation strategies affect?
The lack of transparency in the field of fair machine learning has been acknowledged in literature (Rudin et al., 2020; Wachter et al., 2021; Goethals et al., 2023). Other studies have already specifically criticized the opacity in the effects of the different bias mitigation methods (Krco et al., 2023; Favier et al., 2023; Holstein et al., 2019). What is the impact of each method, not only between sensitive groups but also within each group?
We provide transparency into the operational dynamics of the different bias mitigation methods, by comparing the score distributions after deploying each bias mitigation method with the score distributions of the initial machine learning model. We illustrate the prediction scores from the initial ML model on the x-axis and the scores post-application of various bias mitigation methods on the y-axis. Additionally, we categorize the instances based on their affiliation with either the protected group (represented in dark blue) or the privileged group (represented in light blue). The chart is divided into four quadrants, each depicting the classification of instances as either positive or negative by the initial ML model and by each bias mitigation method. [^8] Note that we use the custom threshold of $0.5$ here to generate the labels, based on the prediction scores. However, as argued by Scantamburlo et al. (2024), this does not take into account the real constraints of decision-making context. We also include a diagonal line that would contain all the instances, if the prediction scores would remain identical. The presented results in Figure 1 are calculated using the Compas dataset. The figures for the other datasets can be found in the Appendix (Figures 3, 4, 5 and 6), with the results being consistent to the results of the Compas dataset.
Insights into the operational dynamics of each bias mitigation method
@@
Figure 1: Score distributions for the Compas dataset. The x-axis represents the prediction scores of the initial ML model, while the y-axis represents the prediction score after applying each bias mitigation method. The second quadrant represents the instances that are ‘upgraded’ by the bias mitigation method (initially predicted as negative, and after using the bias mitigation method predicted as positive), while the fourth quadrant represents the instances that are ‘downgraded’ by the bias mitigation method (initially predicted as positive, and after the bias mitigation method predicted as negative).
In Figure 1(a), we observe the prediction scores of the initial ML model. By default, machine learning models use a classification threshold of $0.5$, categorizing instances above this threshold as positive and those below it as negative. [^9] Note that this is not how it will actually be implemented in deployment. The notion of a decision threshold is not part of the model but is part of the decision logic. This is why it does not make sense to compare bias mitigation methods based on prediction labels generated by the machine learning model, without taking into account the decision-making context.
@@
In Figures 1(b)-1(e), we can assess the relation between the prediction scores of the initial ML model and the prediction scores after applying preprocessing (LFR and DIR) or inprocessing (ADV and MFC) bias mitigation methods. Figure 1(b) reveals that LFR significantly alters the prediction scores, in a seemingly random way. Instances that receive a positive label post-LFR are notably different from those in the initial score ranking. This suggests a substantial transformation in the feature space due to the preprocessing method. We will assess how this affects the performance and fairness metrics in Table 3. As shown in Figures 1(c), 1(d), and 1(e), the methods DIR, ADV, and MFC also introduce significant alterations to the prediction scores, and thus the inherent ranking of instances. However, the altered prediction scores align more closely with the initial prediction scores compared to those obtained using LFR.
@@
Figure 1 also provides us with insights into the fundamental distinction among various bias mitigation methods—specifically, whether they impact the prediction scores or only the prediction labels. Unlike preprocessing and inprocessing methods, the used postprocessing methods do not alter the inherent ranking; instead, they adjust labels based on the prediction scores of the initial machine learning model. Each method implements a unique strategy for label flipping, all with the objective of meeting specified fairness metrics. For instance, ROC (Figure 1(f)) targets the most uncertain instances for label flipping, TO (Figure 1(h)) establishes distinct thresholds for each group, while EOP (Figure 1(g)) determines the ideal quantity of labels to flip within each group and executes these flips randomly. In the scenarios of ROC and TO, group-specific new thresholds are determined (as illustrated in Figures  1(f) and 1(h)). Regarding EOP, label flipping occurs randomly among negatively classified individuals in the protected group and positively classified individuals in the privileged group (as shown in green and red in Figure 1(g)). In this case, the initial ranking impacts the initial labels, but will have no influence on the decision regarding which labels are flipped. [^10] Note that with this postprocessing methods, essentially some reranking does happen, as individuals with a lower prediction score can now receive a positive label, while individuals from the same group with a higher prediction score do not. However, it is not reflected in the prediction scores, so the ranking similarity will be 1.
@@
Our examination provides insights into the specific instances affected by each bias mitigation method, in response to the work of Krco et al. (2023). Take, for instance, the case of ROC: the instances that are flipped from a negative classification to a positive classification, are those instances that were initially classified as negative but that posses the highest prediction score according to the initial machine learning model. We also show in Figure 1 that reranking does happen for some of the methods, but in the next section, we quantify how much.
How similar are the score distributions?
Instead of visualizing how the prediction scores of the initial ML relate to the prediction scores after each bias mitigation method, we can also calculate the overlap between the two rankings by utilizing the Kendall-Tau statistic (Kendall, 1948). The outcomes are presented in Table 2. Table 2 illustrates that the ranking produced by LFR consistently presents the least similarity with the ranking produced by the initial machine learning model. In contrast, the use of DIR leads to a significantly higher degree of overlap in rankings across all datasets. The similarity in rankings when applying ADV and MFC falls into a more moderate category. These observations highlight the varying degrees of impact each method exerts on the prediction scores and thus the ranking of individuals. With postprocessing methods, the ranking remains identical to the ranking of the initial model, resulting in a ranking similarity of 1 (refer to Table 2), and the AUC remains consistent with the initial model (see Table 3). Our analysis reveals that across all datasets, the prediction scores and consequently the intrinsic rankings are substantially modified by every preprocessing and inprocessing bias mitigation method we evaluated. [^11] We see that in one case (MFC for the Law dataset), the scores are even completely shuffled in the opposite way, which shows how random some of the effects can be. While such modifications are not necessarily worrisome, it is crucial to assess whether they enhance or worsen the ranking, and to what extent. It also does not feel fair from the viewpoint of the individual that they are suddenly downweighted by a random intervention.  [^12] So far, we emphasized that only preprocessing and inprocessing methods suffer from this reranking process. However, the postprocessing method EOP also significantly suffers from this arbitrariness, as it will execute random flips within each group until a fairness metric is satisfied. This means that the impacted individuals can be different in each run, and can be individuals with a very high prediction score can be ‘downweighted’ and individuals with a very low prediction score can be ‘uplifted’ (as the prediction scores are not taken into account to determine who should be flipped). Table 2 displays the correlation between each of the bias mitigation methods and the prediction scores of the machine learning model, but we also visualise the correlation between the bias mitigation methods themselves in Figure 2.
@@
Table 2: Table with the similarity between the initial ranking produced by the machine learning model and the ranking produced by the model after using a bias mitigation method, measured by the Kendall-Tau statistic. The values between brackets present the similarity in ranking for the protected group and privileged group respectively.